OpenStudy (anonymous):

How do I solve this problem? Find the exact value of the length of the side of a 45-45-90 right triangle with a hypotenuse of 15 miles.

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

divide by the square root of two

OpenStudy (misty1212):

to find the diagonal if you knew the legs, you would multiply by \(\sqrt2\) to find the legs if you know the diagonal, you divide by \(\sqrt 2\)

OpenStudy (anonymous):

Oh, okay. Thank you! Is this just for special triangles or whatever, orrr...?

OpenStudy (misty1212):

it is for 45 - 45 -90 because it is like the diagonal of a square

OpenStudy (misty1212):



OpenStudy (misty1212):

\[a^2+a^2=h^2\\ 2a^2=h^2\\ \sqrt{2a^2}=h\\ a\sqrt2=h\]

OpenStudy (anonymous):

Okay.

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]